# Dataset Distillation as Pushforward Optimal Quantization

### v2

@@ C.1 Proof of Thm 2
C.2 Convergence of optimal quantization
@@
C.3 Gradient interpretation of CLVQ
Proposition 6 (Differentiability of distortion [34, Prop. 3.1]).
@@ C.3 Gradient interpretation of CLVQ
where $\mathrm{Hull}$ denotes the convex hull. Recalling that the gradient 22 is a $\mu$-expectation over $C_{i}(x)$, the corresponding stochastic Monte Carlo version of the above gradient descent yields the CLVQ algorithm 1:
@@
C.4 Lloyd I algorithm
Lloyd I. This consists of iteratively updating the centroids with the $\mu$-centroids, given by the $\mu$-average of the Voronoi cells. Clearly, if this algorithm converges, then the centroids are equal to the $\mu$-centroids and the grid is stationary. This is more commonly known as the $k$-means clustering algorithm, employed in common numerical software packages such as scikit-learn [37]. Convergence of the Lloyd-I algorithm can be found in e.g. [35].
@@ Appendix D Background on well-posed diffusions
Suppose that the true data distribution on the image space is given by $\mu\in\mathcal{P}(\mathbb{R}^{d})$, assumed to have bounded support. Then, by the Hörmander condition, the law of a random variable $(X_{t})_{t\geq 0}$ evolving under either the VPSDE or the VESDE will admit a density $p_{t}(x)$ for all $t>0$ with respect to the Lebesgue measure, that is smooth with respect to both $x$ and $t$ [18]. Using the following proposition, we have well-definedness of the backward SDE [2, 16].
@@
For the forward SDE 8, assume that there exists some $K>0$ such that
@@
    The solution $X_{t}$ of 8 has a $\mathcal{C}^{1}$ density $p_{t}(x)$ for all $t>0$, and
@@
Then the reverse process $X_{T-t}$ is a solution of the 9, and moreover, the solutions of 9 are unique in law.
Now given that the backwards SDE is indeed a diffusion, the data processing inequality uses the Markov property and states that the divergence after diffusion is less than the divergence before diffusion [23]. This is summarized in [39, Thm. 1].
@@
Denote the initial data distribution by $\mu_{0}=\mu$, and let $\mu_{T}$ be the distribution of a random variable $X_{t}$ satisfying the forward SDE 8 on $[0,T]$. Assume the above assumptions, and let $Y_{t}$ satisfy the backwards SDE 9 on $[0,T]$ with terminal condition $Y_{T}\sim\nu_{T}$. Denote by $\mu_{t}$ and $\nu_{t}$ the marginal distributions at time $t\in[0,T]$ of $X_{t}$ and $Y_{t}$, and assume that $\nu_{T}\ll\mu_{T}$. Then:
@@ Appendix F Proof of main Theorem 1
From [4, Sec. 2.1], the density $p_{t}$ can be written as
@@
We next require the following proposition, which can be thought of as a stochastic version of Gronwall’s inequality. We present a simple version of the even stronger result available in [19].
Proposition 8 ([19, Lem. 3.6]).
@@
Then for processes $X^{x}_{t},X^{y}_{t}$ starting at $x,y$ respectively under 27, it holds for all $t\in(0,T]$ that
@@
Our goal is to “pushforward” the convergence of the Wasserstein distance from the latent space (of optimal quantizations) through the diffusion model (backwards SDE 9) into the image space/manifold. We proceed with the proof.
@@
The main idea is to push a Wasserstein-optimal coupling through the backwards SDE, then using Proposition˜8 and Proposition˜2. First fix a $\delta\in(0,T]$, and let $\operatorname{supp}\mu$ be bounded by $R$. Let $g_{t}=\frac{1}{(2\pi t)^{-d/2}}\exp(-\|x\|^{2}/2t)$ denote the distribution of the Gaussian $\mathcal{N}(0,tI_{d})$ in $d$ dimensions. By the Hörmander condition, the densities of a random variable $X_{t}$ with initial distribution $X_{0}\sim\mu$ undergoing the Brownian motion or Ornstein–Uhlenbeck process exist, and furthermore the forward and backward SDE processes are diffusions [29, 15, 39].
@@
Apply Lemma 1 with $\mu$ for the Brownian motion, and $\mu(e^{t/2}x)$ for the Ornstein–Uhlenbeck process. The corresponding backward SDEs (in forward time) for the Brownian motion and Ornstein–Uhlenbeck process as given by the forward time versions of 9 are
@@ Appendix G Approximate Timings
All times are done on Nvidia A6000 GPUs with 48GB of VRAM. We note that synthesis time as reported in [50, 51] do not include the time required to generate the latent variables, and thus are not sufficiently representative of the end-to-end time required to distill the dataset.
@@ Appendix H Experiment Hyperparameters
We detail the parameters when training the student networks from the distilled data. They are mostly similar to [50].
@@
Variance reduction (heuristic). We note that the variance of the number of cluster assignments can vary significantly, sometimes up to two orders of magnitude, such as in Figure˜1. After normalizing the cluster counts in Step 3 of Algorithm˜2 to give weights $w\in(0,1)$, most weights are very small, and do not contribute much to the neural network training. To reduce this effect, we use the per-centroid weights as follows,
@@
We note that other choices of variance reduction can also be used, and leave a more thorough search to future work.

### v3

@@ C.3 Proof of Proposition 2
For a given quantization $\mathbf{x}$} or quantization level $K$, the quadratic distortion thus satisfies respectively:
C.4 Convergence of optimal quantization
@@
Remark 4.
Convergence to an optimal quantizer is only guaranteed in the case of a log-concave distribution in one dimension (Kieffer, 1982; Liu and Pagès, 2020). In higher dimensions, convergence to a stationary (but not optimal) grid is possible (Pages and Yu, 2016; Pagès, 2015).
C.5 Gradient interpretation of CLVQ
Proposition 6 (Differentiability of distortion Pagès 2015, Prop. 3.1).
@@ C.5 Gradient interpretation of CLVQ
where $\mathrm{Hull}$ denotes the convex hull. Recalling that the gradient 20 is a $\mu$-expectation over $C_{i}(x)$, the corresponding stochastic Monte Carlo version of the above gradient descent yields the CLVQ algorithm 2:
@@
1 Initialize weights ${\mathbf{w}}=(w_{1},...,w_{K})=(1/K,...,1/K)$;
2 while not converged do
3   Sample $X_{i}\sim\mu$;
7 $i\leftarrow i+1$;
8
9 end while
Result: quantization $\nu_{K}=\sum_{k=1}^{K}w_{i}\delta(x_{k}^{*})$
Algorithm 2 CLVQ
C.6 Lloyd I algorithm
Lloyd I. This consists of iteratively updating the centroids with the $\mu$-centroids, given by the $\mu$-average of the Voronoi cells. Clearly, if this algorithm converges, then the centroids are equal to the $\mu$-centroids and the grid is stationary. This is more commonly known as the $k$-means clustering algorithm, employed in common numerical software packages such as scikit-learn (Pedregosa et al., 2011). Convergence of the Lloyd-I algorithm can be found in e.g. (Pages and Yu, 2016).
@@ Appendix D Background on well-posed diffusions
Suppose that the true data distribution on the image space is given by $\mu\in\mathcal{P}(\mathbb{R}^{d})$, assumed to have bounded support. Then, by the Hörmander condition, the law of a random variable $(X_{t})_{t\geq 0}$ evolving under either the VPSDE or the VESDE will admit a density $p_{t}(x)$ for all $t>0$ with respect to the Lebesgue measure, that is smooth with respect to both $x$ and $t$ (Hörmander, 1967). Using the following proposition, we have well-definedness of the backward SDE (Anderson, 1982; Haussmann and Pardoux, 1986).
@@
For the forward SDE 5, assume that there exists some $K>0$ such that
@@
    The solution $X_{t}$ of 5 has a $\mathcal{C}^{1}$ density $p_{t}(x)$ for all $t>0$, and
@@
Then the reverse process $X_{T-t}$ is a solution of the 6, and moreover, the solutions of 6 are unique in law.
Now given that the backwards SDE is indeed a diffusion, the data processing inequality uses the Markov property and states that the divergence after diffusion is less than the divergence before diffusion (Liese and Vajda, 2006). This is summarized in (Pidstrigach, 2022, Thm. 1).
@@
Denote the initial data distribution by $\mu_{0}=\mu$, and let $\mu_{T}$ be the distribution of a random variable $X_{t}$ satisfying the forward SDE 5 on $[0,T]$. Assume the above assumptions, and let $Y_{t}$ satisfy the backwards SDE 6 on $[0,T]$ with terminal condition $Y_{T}\sim\nu_{T}$. Denote by $\mu_{t}$ and $\nu_{t}$ the marginal distributions at time $t\in[0,T]$ of $X_{t}$ and $Y_{t}$, and assume that $\nu_{T}\ll\mu_{T}$. Then:
@@ Appendix F Proof of main Theorem 1
From (Bardet et al., 2018, Sec. 2.1), the density $p_{t}$ can be written as
@@
We next require the following proposition, which can be thought of as a stochastic version of Gronwall’s inequality. We present a simple version of the even stronger result available in Hudde et al. (2021).
Proposition 8 (Hudde et al. 2021, Lem. 3.6).
@@
Then for processes $X^{x}_{t},X^{y}_{t}$ starting at $x,y$ respectively under 25, it holds for all $t\in(0,T]$ that
@@
Our goal is to “pushforward” the convergence of the Wasserstein distance from the latent space (of optimal quantizations) through the diffusion model (backwards SDE 6) into the image space/manifold. We proceed with the proof.
@@
The main idea is to push a Wasserstein-optimal coupling through the backwards SDE, then using Proposition˜8 and Proposition˜2. First fix a $\delta\in(0,T]$, and let $\operatorname{supp}\mu$ be bounded by $R$. Let $g_{t}=\frac{1}{(2\pi t)^{-d/2}}\exp(-\|x\|^{2}/2t)$ denote the distribution of the Gaussian $\mathcal{N}(0,tI_{d})$ in $d$ dimensions. By the Hörmander condition, the densities of a random variable $X_{t}$ with initial distribution $X_{0}\sim\mu$ undergoing the Brownian motion or Ornstein–Uhlenbeck process exist, and furthermore the forward and backward SDE processes are diffusions (Malliavin, 1978; Hairer, 2011; Pidstrigach, 2022).
@@
Apply Lemma 1 with $\mu$ for the Brownian motion, and $\mu(e^{t/2}x)$ for the Ornstein–Uhlenbeck process. The corresponding backward SDEs (in forward time) for the Brownian motion and Ornstein–Uhlenbeck process as given by the forward time versions of 6 are
@@ Appendix G Approximate Timings
All times are done on Nvidia A6000 GPUs with 48GB of VRAM. We note that synthesis time as reported in Su et al. (2024); Sun et al. (2024) do not include the time required to generate the latent variables, and thus are not sufficiently representative of the end-to-end time required to distill the dataset.
@@ Appendix H Experiment Hyperparameters
We detail the parameters when training the student networks from the distilled data. They are mostly similar to Su et al. (2024).
For consistency and a more direct comparison with previous methods, we use the pre-trained PyTorch ResNet-18 model to compute the soft labels, using the same protocol as Su et al. (2024). After computing the soft labels using the pre-trained ResNet-18 model, we train new ResNet-18, ResNet-50 and ResNet-101 models to match the soft labels. The data augmentation is also identical, so that the only differences are the addition of the weights to the training objective and some minor hyperparameter tuning for the new objective. The latent diffusion model chosen for latent generation and image synthesis is the publicly available pre-trained Stable Diffusion V1.5 model, the same as D4M.
@@
Variance reduction (heuristic). We note that the variance of the number of cluster assignments can vary significantly, sometimes up to two orders of magnitude, such as in Figure˜2. After normalizing the cluster counts in Step 3 of Algorithm˜1 to give weights $w\in(0,1)$, most weights are very small, and do not contribute much to the neural network training. To reduce this effect, we use the per-centroid weights as follows,
@@
Appendix I Ablation on weighting
To test the effect of the heuristic weights used in 34, we consider the setting of the DiT backbone Section˜4.2. We consider three different weightings within the training in Step 4:
  1. 1.
    Constant weights $w\equiv 1$. This is equivalent to D4M.
  2. 2.
    The heuristic weighting strategy 34.
  3. 3.
    The direct cluster weights, given by $w_{k}\propto v_{k}^{-1}$, where $v_{k}$ are the cluster counts from the $k$-means, normalized such that the sum over each class is 1.
We use the same cluster images and only differ the weights. Table˜7 details the test accuracy for IPC 10 on ImageNet-1K with various weights, applied with a ResNet-18 teacher and ResNet-18 student architectures. We observe that the heuristic weighting strategy outperforms D4M across different learning rates. Moreover, the direct cluster weights are able to obtain higher test accuracy, albeit being more sensitive to the choice of learning rate.
Table 7: Ablation across different weighting strategies when training a new student network with soft labels. We observe that the heuristic weight used in the main experiments always outperforms D4M. The direct cluster weighting is more sensitive to learning rate, but eventually outperforms the heuristic weighting for higher learning rates.
Learning rate 1e-3 2e-3 5e-3 1e-2
No weighting (D4M) 52.1 52.5 51.5 52.2
Heuristic weight 34 52.3 52.6 53.2 53.6
Direct cluster weighting 44.1 50.3 54.6 55.6